# LV-Eval: A Balanced Long-Context Benchmark with 5 Length Levels Up to 256K

### v2

@@ Abstract
State-of-the-art large language models (LLMs) are now claiming remarkable supported context lengths of $256k$ or even more. In contrast, the average context lengths of mainstream benchmarks are insufficient ($5k$-$21k$), and they suffer from potential knowledge leakage and inaccurate metrics, resulting in biased evaluation. This paper introduces LV-Eval, a challenging long-context benchmark with five length levels ($16k$, $32k$, $64k$, $128k$, and $256k$) reaching up to $256k$ words. LV-Eval features two main tasks, single-hop QA and multi-hop QA, comprising 11 bilingual datasets. The design of LV-Eval has incorporated three key techniques, namely confusing facts insertion (CFI), keyword and phrase replacement (KPR), and keyword-recall-based metric design. The advantages of LV-Eval include controllable evaluation across context lengths, challenging test instances with confusing facts, mitigated knowledge leakage, and more objective evaluation. We evaluate 15 LLMs on LV-Eval and conduct ablation studies on the benchmarking techniques. The results reveal that: (i) Moonshot-v1 and recent large-scale open-source models, such as Qwen-2.5-72B and Llama-3.1-70B, achieve the highest performance on LV-Eval, particularly at lengths below $64k$. (ii) Models exhibit distinct score trends. For example, GLM-4-9B-128k, Yi-6B-200k, and Llama3-8B-1M exhibit a relatively gentle degradation of performance, but their absolute performances may not necessarily be higher than those of LLMs with shorter context lengths. (iii) LLMs’ performances can significantly degrade in the presence of confusing information, especially in the pressure test of “needle in a haystack”. (iv) Issues related to knowledge leakage and inaccurate metrics introduce bias in evaluation, and these concerns are alleviated in LV-Eval. All datasets and evaluation codes are released at: https://github.com/infinigence/LVEval.
@@ 1 Introduction
BAMBOO Dong et al. (2023) 10 5,067 229/14,858 $4k$,$16k$ en+zh
LongBench Bai et al. (2023b) 21 9,486 128/71,954 00-$4k$,$4k$-$8k$,$8k$+ en+zh
LV-Eval 11 102,380 11,896/387,406 $16k$,$32k$,$64k$,$128k$,$256k$ en+zh
@@ 2 Related Work
[…]
@@ 3 LV-Eval Benchmark
Datasets in LV-Eval are constructed with existing public datasets as the source, except for factrecall-en and factrecall-zh, which are constructed using the data from PG19 (Rae et al., 2019) dataset and Journey to the West book. Each dataset consists of five subsets of different lengths: $16k$, $32k$, $64k$, $128k$, and $256k$. All five subsets share the same question-answer (QA) pairs, meaning there are five contexts of varying lengths for each QA pair. This allows for a controllable evaluation of models’ long-context ability when testing the same set of questions with different context lengths. In total, LV-Eval comprises 1,729 QA pairs and 1,729$\times$5 = 8,645 synthetic contexts.
@@ 4.1 Compare LLMs on LV-Eval
Figure 4: Overall results on different length levels and types of datasets. (a) Average scores across all datasets of 15 LLMs at 5 length levels. The circle markers represent open-source models of 6-9B; The square markers represent larger open-source models around 70B; The triangle markers represent commercial APIs. Note that we only evaluate Moonshot-v1-128k on $16k$ and $32k$ length levels due to the high cost. (b) Average scores across all length levels of 14 LLMs on 5 types of datasets. “CQA” refers to QA datasets with CFI.
@@ Acknowledgments
We thank the insightful discussion with Kai Chen and Songyang Zhang from Shanghai Artificial Intelligence Laboratory. Thanks to all of the annotators who contribute to the project. We thank Hanling Zhang for helping run the experiment for Moonshot.
@@ Appendix A Detailed Construction of QA Pairs
[…]
@@ Appendix C Detailed Evaluation Results
Figure A8: Average scores across all length levels of 14 LLMs on single-hop QA datasets.
Figure A9: Average scores across all length levels of 14 LLMs on multi-hop QA datasets.

### v3

@@ Abstract
State-of-the-art large language models (LLMs) are now claiming remarkable supported context lengths of $256k$ or even more. In contrast, the average context lengths of mainstream benchmarks are insufficient (5k5k-$21k$), and they suffer from potential knowledge leakage and inaccurate metrics, resulting in biased evaluation. This paper introduces LV-Eval, a challenging long-context benchmark with five length levels ($16k$, $32k$, $64k$, $128k$, and $256k$) reaching up to $256k$ words. LV-Eval features two main tasks, single-hop QA and multi-hop QA, comprising 11 bilingual datasets. The design of LV-Eval has incorporated three key techniques, namely confusing facts insertion (CFI), keyword and phrase replacement (KPR), and keyword-recall-based metric design. The advantages of LV-Eval include controllable evaluation across context lengths, challenging test instances with confusing facts, mitigated knowledge leakage, and more objective evaluation. We evaluate 15 LLMs on LV-Eval and conduct ablation studies on the benchmarking techniques. The results reveal that: (i) Moonshot-v1 and recent large-scale open-source models, such as Qwen-2.5-72B and Llama-3.1-70B, achieve the highest performance on LV-Eval, particularly at lengths below $64k$. (ii) Models exhibit distinct score trends. For example, GLM-4-9B-128k, Yi-6B-200k, and Llama3-8B-1M exhibit a relatively gentle degradation of performance, but their absolute performances may not necessarily be higher than those of LLMs with shorter context lengths. (iii) LLMs’ performances can significantly degrade in the presence of confusing information, especially in the pressure test of “needle in a haystack”. (iv) Issues related to knowledge leakage and inaccurate metrics introduce bias in evaluation, and these concerns are alleviated in LV-Eval. All datasets and evaluation codes are released at: https://github.com/infinigence/LVEval.
@@ 1 Introduction
BAMBOO Dong et al. (2023) 10 5,067 229/14,858 4k4k,$16k$ en+zh
LongBench Bai et al. (2023b) 21 9,486 128/71,954 0-4k4k,4k4k-8k8k,8k8k+ en+zh
LV-Eval 11 102,380 11,896/387,406 16k16k,32k32k,64k64k,128k128k,$256k$ en+zh
@@ 2 Related Work
Table 1 provides a summary of existing long-context benchmarks, including ZeroScrolls (Shaham et al., 2023), LooGLE (Li et al., 2023b), L-Eval (An et al., 2023), BAMBOO (Dong et al., 2023), and LongBench (Bai et al., 2023b). ZeroScrolls, LooGLE, and L-Eval are monolingual benchmarks without explicit length level partition. Their average word counts are $\sim$$14k$, $\sim$$21k$ and $\sim$$13.5k$, respectively. In order to evaluate the model’s capability across various context lengths, BAMBOO and LongBench have designed various length levels. However, the word counts ($\sim$$5k$, $\sim$$9.5k$) of the contexts in these two benchmarks are notably smaller than the supported context length of state-of-the-art long-context models, making them unsuitable for evaluating the claimed extremely long-context understanding ability. In contrast, LV-Eval contains five length levels, up to $256k$ words, each with the same set of QA pairs for controllable evaluation.
@@ 3 LV-Eval Benchmark
[…]
@@ 4.1 Compare LLMs on LV-Eval
Figure 4: Overall results on different length levels and types of datasets. (a) Average scores across all datasets of 15 LLMs at 5 length levels. The circle markers represent open-source models of 6-9B; The square markers represent larger open-source models around 70B; The triangle markers represent commercial APIs. Note that we only evaluate Moonshot-v1-128k on $16k$ and $32k$ length levels due to the high cost. (b) Average scores across all length levels of 15 LLMs on 5 types of datasets. “CQA” refers to QA datasets with CFI.
@@ Acknowledgments
We thank the insightful discussion with Kai Chen and Songyang Zhang from Shanghai Artificial Intelligence Laboratory. Thanks to all of the annotators who contribute to the project. We thank Hanling Zhang for helping run the experiment for Moonshot. We thank Qin Cheng from Yale University for presenting our poster at the Conference on Language Modeling (COLM 2025).
@@ Appendix A Detailed Construction of QA Pairs
  • cmrc-mixup is derived from the CMRC 2018 Public Datasets, designed for Chinese machine reading comprehension. It contains $\sim$$20k$ questions annotated on Wikipedia documents by human experts. We manually pick 200 QA pairs and their corresponding documents as supporting QA pairs and documents.
@@ Appendix C Detailed Evaluation Results
Figure A8: Average scores across all length levels of 15 LLMs on single-hop QA datasets.
Figure A9: Average scores across all length levels of 15 LLMs on multi-hop QA datasets.